# Qincheng Lu School of Computer Science

### v1

@@ Abstract
Massive MIMO (multiple-input multiple-output) detection is an important topic in wireless communication and various machine learning based methods have been developed recently for this task. Expectation propagation (EP) and its variants are widely used for MIMO detection and have achieved the best performance. However, EP-based solvers fail to capture the correlation between unknown variables, leading to loss of information, and in addition, they are computationally expensive. In this paper, we show that the real-valued system can be modeled as spectral signal convolution on graph, through which the correlation between unknown variables can be captured. Based on this analysis, we propose graph convolution-enhanced expectation propagation (GCEPNet), a graph convolution-enhanced EP detector. GCEPNet incorporates data-dependent attention scores into Chebyshev polynomial for powerful graph convolution with better generalization capacity. It enables a better estimation of the cavity distribution for EP and empirically achieves the state-of-the-art (SOTA) MIMO detection performance with much faster inference speed. To our knowledge, we are the first to shed light on the connection between the system model and graph convolution, and the first to design the data-dependent attention scores for graph convolution.
@@
massive MIMO detection, machine learning, graph convolution, expectation propagation
@@ I Introduction
Massive MIMO (multiple-input multiple-output) is a wireless communication technology that uses a very large number of antennas at the base station to significantly increase spectral efficiency and capacity [1]. One of the outstanding challenges in massive MIMO is to efficiently detect the transmitted signal [2]. Finding the optimal detector can be modeled as solving an integer least square problem, which is NP-hard [3]. As the complexity of the traditional methods for near-optimal solutions still grows prohibitively as the number of users increases, especially for large-scale problem, methods inspired by deep learning have emerged as promising solutions to meet the high throughput and low latency requirements [4]. Achieving a comparable detection performance to the traditional methods, the learning-based methods shift the significant part of computational cost to an offline training phase and significantly shorten run time in the online detection (inference) phase [5].
@@
For improving the performance of conventional detectors [6, 7, 8, 9, 10, 11], various machine learning methods for MIMO detection have been proposed [5], including pure learning-based methods [12, 13], and the methods which incorporate learnable parameters [14, 2, 15, 16] or learnable modules [17, 18, 19, 20] into some conventional detectors. Among them, graph neural network (GNN) aided massage passing (MP) detectors [17, 18, 19, 20] stands out with better performance. In GNN aided MP detectors, GNN is used to improve the approximation of posterior distribution in various MP detectors, including Expectation Propagation (EP) [10], Approximate Message Passing (AMP) [11] and Bayesian parallel interference cancellation (BPIC) [21]. The GNN aided EP framework (GEPNet) [17] and its variant IEP-GNN [20] (to be elaborated later) achieve the best performance among well-known conventional sub-optimal detectors and those which apply machine learning techniques.
@@
Here is a brief history about the development of incorporating GNNs into MP. GEPNet is the first one, but the computation of GEPNet involves matrix inversion due to the EP iterations. As it is inefficient, later efforts were made to leverage other MP methods in GEPNet. For example, GPICNet [18] substitutes the EP module with BPIC, and AMP-GNN [19] replaces it by AMP. Although both changes reduce complexity, they degrade GEPNet’s performance. Very recently, IEP-GNN [20] was proposed to improve the computation of local posterior information in the EP iteration, which enhances the performance of GEPNet while maintaining the same computational cost. However, we have not seen any work that improves efficiency of the GNN module in GEPNet. This paper aims to address the inefficiency issue with the GNN module in GEPNet with no degradation and possibly even enhancement of the performance of GEPNet. The proposed method can be seamlessly integrated with IEP-GNN.
@@
Meanwhile, the theoretical justification, expressivity, and efficiency of the currently widely adopted GNNs framework for MIMO detection are under-explored. Existing GNNs for MIMO are intuitively based on the pair-wise Markov Random Field (MRF) factorization of the posterior [12, 17, 18], where a spatial GNN is utilized to parameterize and learn the representation of the variable and factor nodes in the MRF model. However, the formulation of these GNNs inherits from works that combine deep learning with MRF [22, 12], but less motivated by the MIMO detection problem. In addition, spatial GNNs originate as the low order approximation of the graph convolution for graph signal processing [23]. To enhance the expressivity of the GNN for MIMO detection, we investigate the connection between the system model and the graph convolution, and propose a novel graph convolution-enhanced EP detector (GCEPNet)  [^1] Code available at: https://anonymous.4open.science/r/GCEPNet-F118/, which achieves state-of-the-art detection performance with fewer training parameters and lower inference complexity.
@@ II-A Notation
We use bold capital and small letters for matrices and vectors, respectively. The $(i,j)$ entry of $\mathbf{A}$ is denoted by $\mathbf{A}(i,j)$ or $a_{ij}$ the $i^{th}$ row of $\mathbf{A}$ is denoted by $\mathbf{A}_{i,:}$, and the $i^{th}$ entry of $\mathbf{s}$ is denoted by $\mathbf{s}(i)$ or $s_{i}$. We denote ${\mathbf{1}}=[1,\ldots,1]^{\mkern-1.5mu\mathsf{T}}$. When a real-valued random vector ${\mathbf{x}}$ follows the normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix ${\boldsymbol{\Sigma}}$, we write ${\mathbf{x}}\sim\mathcal{N}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ and use $N({\mathbf{x}}\!:\!\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ to denote its density function.
@@ VI-B Architecture of GCEPNet
where the $i$-th row of $\mathbf{S}^{(0)}$ is a graph signal of the node $i$ corresponding to $x_{i}$.  [^2] A remark about (24): Since the noise vector $\mathbf{n}\in\mathcal{N}({\mathbf{0}},\sigma_{n}^{2}\mathbf{I})$ in equation (23) is unknown, so we use $\sigma_{n}{\mathbf{1}}$ to represent it. To mitigate this approximation error we separate $\alpha\mathbf{H}^{\mkern-1.5mu\mathsf{T}}{\mathbf{y}}$ and $\alpha\sigma_{n}\mathbf{H}^{\mkern-1.5mu\mathsf{T}}\mathbf{1}$ to define $\mathbf{S}^{(0)}$ in equation (24) so that the two vectors can be scaled separately later (see equation (26)). The key operation in GCEPNet is that we use the graph convolution in equation (23) to recover useful information about the unknown ${\mathbf{x}}$ from the graph signal defined in (24), by the following approximated Chebyshev polynomials with order $M$,
@@
At $t^{th}$ EP iteration, based on the cavity density of each $x_{i}$ in (18), we construct $\mathbf{a}^{(t)}=[a_{0}^{(t)},\cdots,a_{2N_{t}}^{(t)}]^{\mkern-1.5mu\mathsf{T}}$, $\mathbf{b}^{(t)}=[(b_{0}^{(t)})^{2},\cdots,(b_{2N_{t}}^{(t)})^{2}]^{\mkern-1.5mu\mathsf{T}}$, and denote $\mathbf{P}^{(t)}=[\mathbf{a}^{(t)},\mathbf{b}^{(t)}]$, where $\mathbf{P}^{(t)}$ contains the mean and variance information for $x_{i}$. GCEPNet intends to improve the standard cavity density by incorporating pair-wise information learned via the graph convolution on ${\cal G}_{\text{MIMO}}$ efficiently. In Figure 1, we present the overall structure of GCEPNet, where the graph convolution and the GRU module at the $t^{th}$ EP iteration are
@@
$\hat{\mathbf{S}}^{(t)}=\left(\sum_{m=0}^{M}\text{ATT}(\bar{\mathbf{S}}^{(0)})_{m}T_{m}(\mathbf{L}_{\text{MIMO}})\right)\widetilde{\mathbf{S}}^{(t)},$ (28)
@@
$\mathbf{P}_{\mathcal{G}}^{(t)}=R\left(\text{MLP}_{2}(\bar{\mathbf{S}}^{(t+1)})\right).$ (30)
At the initial EP iteration, in (26), trainable parameters $\mathbf{W}_{0}\in{\mathbb{R}}^{2\times N_{u}}$ and $\mathbf{B}_{0}\in{\mathbb{R}}^{2N_{t}\times N_{u}}$ transform the initial 2-dimensional graph signal $\mathbf{S}^{(0)}$ to a hidden representation $\bar{\mathbf{S}}^{(0)}$ with size $N_{u}$. We use MLP for a Multi-Layer Perceptron (MLP). The graph convolution in (28) is applied to the input graph signal $\widetilde{\mathbf{S}}^{(t)}$ to compute the output signal $\hat{\mathbf{S}}^{(t)}$. The attention-based network $\text{ATT}(\bar{\mathbf{S}}^{(0)})$ learns an $M+1$ dimensional vector containing data-dependent Chebyshev coefficients $\text{ATT}(\bar{\mathbf{S}}^{(0)})_{m}$ for $m=0,1,\ldots,M$. The gated recurrent unit (GRU) [27] in (29) provides a gating mechanism to integrate the output signal $\hat{\mathbf{S}}^{(t)}$ and the cavity distribution stored in $\mathbf{P}^{(t)}$. The output of the GRU will be the input of the GNN module at the next EP iteration. The readout module $R$ in (30) computes the mean and variance estimation of each $x_{i}$. Two GNN layers perform (27)-(29) two times. More explanations are given in the following two subsections.
@@ VI-C Data-dependent Graph Convolution Coefficients
In equation (5), ChebNet uses conventional Chebyshev polynomial as graph filter, which uses fixed coefficients $w_{m}$. However, this architecture is less powerful and cannot perform well on testing data. Thus, in (28), we proposed to replace the fixed $w_{m}$ by data-dependent learnable attention scores $\text{ATT}(\bar{\mathbf{S}}^{(0)})_{m}$. Training $\text{ATT}(\bar{\mathbf{S}}^{(0)})$, which is a function of an instance $\{{\mathbf{y}},\mathbf{H},\sigma_{n}\}$, can improve the generalization ability of the graph convolution module. $\text{ATT}(\bar{\mathbf{S}}^{(0)})$ first computes attention score $\alpha_{i}$ between the default coefficient vector ${\mathbf{1}}$ with size $M+1$ and each row of the key matrix $\mathbf{K}\in{\mathbb{R}}^{2N_{t}\times(M+1)}$ as follows
$\alpha_{i}=\frac{\exp(\mathbf{K}_{i,:}{\mathbf{1}})}{\sum_{j}\exp(\mathbf{K}_{j,:}{\mathbf{1}})},\;\text{where }\mathbf{K}=\text{MLP}_{3}(\bar{\mathbf{S}}^{(0)}).$ (31)
The $m$-th column of the key matrix $\mathbf{K}$ is an embedding for the $m$-th coefficient. The output of $\text{ATT}(\bar{\mathbf{S}}^{(0)})$ is a weighted average of each dimension of the coefficient embedding in rows of $\mathbf{K}$, where the attention score $\alpha_{i}$ serves as the normalized weight.
$\text{ATT}(\bar{\mathbf{S}}^{(0)})=\left(\sum_{i}\alpha_{i}\mathbf{K}_{i,:}\right)^{\mkern-1.5mu\mathsf{T}}.$ (32)
To our best knowledge, we are the first to introduce data-dependent attention scores in polynomial filters to increase the expressive power of graph neural networks.
@@ VII Computational Complexity Analysis
For GCEPNet, the $\text{MLP}_{1}$ in (27) has an input and an output layer of size $N_{u}$, and the $\text{MLP}_{3}$ in (31) has an input of size $N_{u}$ and an output of size $M$, then the total cost for the graph convolution in (21), (27), (28), (31) and (32) is $\mathcal{C}_{2}=N(N_{u}N_{h_{1}}+N_{h_{1}}N_{h_{2}}+N_{h_{2}}N_{u}+M)+N^{2}N_{u}M$. When $M<433$, $\mathcal{C}_{2}<\mathcal{C}_{1}$. With a moderate $M$, $\mathcal{C}_{2}$ is much smaller than $\mathcal{C}_{1}$. In practice, we use $M=3$. Figure 2 demonstrates that, comparing with EP, GEPNet  [^3] Following the implementations in [17]: https://github.com/GNN-based-MIMO-Detection/GNN-based-MIMO-Detection does not scale with the problem size as a result of the inefficient GNN aggregation, while GCEPNet effectively resolves the bottleneck with the newly proposed graph convolution.
Figure 2: The total inference running time (in seconds) for $5\times 10^{4}$ samples with $N_{t}=N_{r}$ and the same number of EP iterations ($T=10$). Both GCEPNet and GEPNet use 2 GNN layers in each EP iteration. The implementation of GEPNet is from [17]. All methods are under CUDA acceleration.
@@ VIII Numerical Results
With the average sample loss, the training uses 1000 epochs to optimize parameters in each network. There are 100 iterations per epoch and each iteration contains 100 training samples. Every training sample has a $\text{SNR}_{\text{dB}}$ uniformly sampled from $[20,60]$. In each epoch, the network is tested on a validation set containing 2000 samples for each $\text{SNR}_{\text{dB}}$ from $\{20,21,\cdots,60\}$. The training uses the Adam optimizer with a learning rate of $10^{-3}$, which is adjusted very 100 epochs according to the validation loss by the ReduceLROnPlateau scheduler in PyTorch. The network with the best mean SER on the validation set is selected for the testing stage. All networks are trained for each problem size and QAM configuration.
The testing uses $5\times 10^{4}$ samples for each $\text{SNR}_{\text{dB}}$. Figure 3 compares the average SER of GCEPNet, GEPNet [17] and EP [10] for $N_{t}=N_{r}=16$ and $N_{t}=N_{r}=32$ with 64-QAM. Both GCEPNet and GEPNet enhance the performance of EP by incorporating correlation information into the standard cavity estimation. While GCEPNet is significantly more efficient than GEPNet, it consistently outperforms GEPNet as well. Given that GEPNet achieves the best performance among existing sub-optimal detectors [18], GCEPNet is established as the new state-of-the-art (SOTA) method.

### v2

@@ Abstract
Massive MIMO (multiple-input multiple-output) detection is an important topic in wireless communication and various machine learning based methods have been developed recently for this task. Expectation Propagation (EP) and its variants are widely used for MIMO detection and have achieved the best performance. However, EP-based solvers fail to capture the correlation between unknown variables, leading to a loss of information, and in addition, they are computationally expensive. In this paper, we show that the real-valued system can be modeled as spectral signal convolution on graph, through which the correlation between unknown variables can be captured. Based on such analysis, we propose graph convolution-enhanced expectation propagation (GCEPNet). GCEPNet incorporates data-dependent attention scores into Chebyshev polynomial for powerful graph convolution with better generalization capacity. It enables a better estimation of the cavity distribution for EP and empirically achieves the state-of-the-art (SOTA) MIMO detection performance with much faster inference speed. To our knowledge, we are the first to shed light on the connection between the system model and graph convolution, and the first to design the data-dependent coefficients for graph convolution.
@@
massive MIMO detection, machine learning for communication, graph convolution, expectation propagation
@@ I Introduction
Massive MIMO (multiple-input multiple-output) is a wireless communication technology that uses a very large number of antennas at the base station to significantly increase spectral efficiency and capacity [1]. One of the outstanding challenges in massive MIMO is to efficiently detect the transmitted signal. Finding the optimal detector can be modeled as solving an integer least square problem, which is NP-hard [2]. As the complexity of the traditional methods for near-optimal solutions still grows prohibitively as the number of users increases, especially for large-scale problem, methods inspired by deep learning have emerged as promising solutions to meet the high throughput and low latency requirements [3]. Achieving a comparable detection performance to the traditional methods, the learning-based methods shift the significant part of computational cost to an offline training phase and significantly shorten run time in the online detection (inference) phase [4].
@@
For improving the performance of conventional detectors [5, 6, 7, 8, 9, 10], various machine learning methods for MIMO detection have been proposed [4], including pure learning-based methods [11, 12], and the methods which incorporate learnable parameters [13, 14, 15] or learnable modules [16, 17, 18, 19] into some conventional detectors. Among them, graph neural network (GNN) aided massage passing (MP) detectors [16, 17, 18, 19] stands out with better performance. In GNN aided MP detectors, GNN is used to improve the approximation of posterior distribution in various MP detectors, including Expectation Propagation (EP) [9], Approximate Message Passing (AMP) [10] and Bayesian parallel interference cancellation (BPIC) [20]. The GNN aided EP framework (GEPNet) [16] and its variant IEP-GNN [19] (to be elaborated later) achieve the best performance among well-known conventional sub-optimal detectors and those which apply machine learning techniques.
@@
Here is a brief history about the incorporation of GNNs into MP. GEPNet is the pioneering work, but its computation involves matrix inversion due to the EP iterations [16]. As it is inefficient, later efforts were made to leverage other MP methods in GEPNet. For example, GPICNet [17] substitutes the EP module with BPIC, and AMP-GNN [18] replaces it by AMP. Although both changes reduce complexity, they degrade GEPNet’s performance. Very recently, IEP-GNN [19] was proposed to improve the computation of local posterior information in the EP iteration, which enhances the performance of GEPNet while maintaining the same computational cost. However, to our best knowledge, there is no existing work that aims to improve the efficiency of the GNN module in GEPNet. In this paper, we focus on addressing the inefficiency issue of the GNN module in GEPNet, with no performance degradation and sometimes even performance enhancement. The proposed method can be seamlessly integrated with IEP-GNN.
@@
Meanwhile, the theoretical justification, expressivity, and efficiency of the currently widely adopted GNNs framework for MIMO detection are under-explored. For example, existing GNNs for MIMO are intuitively based on the pair-wise Markov Random Field (MRF) factorization of the posterior [11, 16, 17], where a spatial GNN is utilized to parameterize and learn the representation of the variable and factor nodes in the MRF model. However, the formulation of these GNNs inherits from works that combine deep learning with MRF [11], but less motivated by the unique property of MIMO detection problem. In addition, spatial GNNs originate as the low order approximation of the graph convolution [21], which loses high-order topology information. To enhance the expressivity of the MIMO-specific GNNs, we investigate the connection between the system model and the graph convolution, and propose a novel graph convolution-enhanced EP detector (GCEPNet)  [^1] Code available at: https://github.com/wzzlcss/GCEPNet, which achieves state-of-the-art detection performance with fewer training parameters and lower inference complexity.
@@ II-A Notation
We use bold capital and small letters for matrices and vectors, respectively. The $(i,j)$ entry of $\mathbf{A}$ is denoted by $\mathbf{A}(i,j)$ or $a_{ij}$, the $i^{th}$ row of $\mathbf{A}$ is denoted by $\mathbf{A}_{i,:}$, and the $i^{th}$ entry of $\mathbf{s}$ is denoted by $\mathbf{s}(i)$ or $s_{i}$. We denote ${\mathbf{1}}=[1,\ldots,1]^{\mkern-1.5mu\mathsf{T}}$. When a real-valued random vector ${\mathbf{x}}$ follows the normal distribution with mean $\boldsymbol{\mu}$ and covariance matrix ${\boldsymbol{\Sigma}}$, we write ${\mathbf{x}}\sim\mathcal{N}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ and use $N({\mathbf{x}}\!:\!\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ to denote its density function.
@@ VI-B Architecture of GCEPNet
where $\alpha\sigma_{n}\mathbf{H}^{\mkern-1.5mu\mathsf{T}}\mathbf{1}$ is used to approximate the unknown $\alpha\mathbf{H}^{\mkern-1.5mu\mathsf{T}}\mathbf{n}$ in (23), and the $i$-th row of $\mathbf{S}^{(0)}$ is a graph signal of the node $i$ corresponding to $x_{i}$. The key operation in GCEPNet is that we use the graph convolution in equation (23) to recover useful information about the unknown ${\mathbf{x}}$ from the graph signal defined in (24), by the following approximated Chebyshev polynomials with order $M$,
@@
At $t^{th}$ EP iteration, based on the cavity density of each $x_{i}$ in (18), we construct $\mathbf{a}^{(t)}=[a_{0}^{(t)},\cdots,a_{2N_{t}}^{(t)}]^{\mkern-1.5mu\mathsf{T}}$, $\mathbf{b}^{(t)}=[(b_{0}^{(t)})^{2},\cdots,(b_{2N_{t}}^{(t)})^{2}]^{\mkern-1.5mu\mathsf{T}}$, and denote $\mathbf{P}^{(t)}=[\mathbf{a}^{(t)},\mathbf{b}^{(t)}]$, where $\mathbf{P}^{(t)}$ contains the mean and variance information for $x_{i}$. GCEPNet intends to improve the standard cavity density by incorporating pair-wise information learned via the graph convolution on ${\cal G}_{\text{MIMO}}$ efficiently. In Figure 1, we present the overall structure of GCEPNet. For $t=0\dots T$,
@@
$\hat{\mathbf{S}}^{(t)}=\textstyle\sum_{m=0}^{M}\text{ATT}(\mathbf{S}^{(0)})_{m}T_{m}(\mathbf{L}_{\text{MIMO}})\widetilde{\mathbf{S}}^{(t)},$ (28)
@@
$\mathbf{P}_{\mathcal{G}}^{(t)}=R(\text{MLP}_{2}(\bar{\mathbf{S}}^{(t+1)})).$ (30)
Equation (26) is to mitigate the error caused by replacing $\alpha\mathbf{H}^{\mkern-1.5mu\mathsf{T}}\mathbf{n}$ with $\alpha\sigma_{n}\mathbf{H}^{\mkern-1.5mu\mathsf{T}}\mathbf{1}$ in (24). At the initial EP iteration, in (26), trainable parameters $\mathbf{W}_{0}\in{\mathbb{R}}^{2\times N_{u}}$ and $\mathbf{B}_{0}\in{\mathbb{R}}^{2N_{t}\times N_{u}}$ transform the initial 2-dimensional graph signal $\mathbf{S}^{(0)}$ to a hidden representation $\bar{\mathbf{S}}^{(0)}$ with size $N_{u}$. We use MLP for a Multi-Layer Perceptron (MLP). The graph convolution in (28) is applied to the input graph signal $\widetilde{\mathbf{S}}^{(t)}$ to compute the output signal $\hat{\mathbf{S}}^{(t)}$. The attention-based network $\text{ATT}(\mathbf{S}^{(0)})$ learns an $M+1$ dimensional vector containing data-dependent Chebyshev coefficients $\text{ATT}(\mathbf{S}^{(0)})_{m}$ for $m=0,1,\ldots,M$. The gated recurrent unit (GRU) [25] in (29) provides a gating mechanism to integrate the output signal $\hat{\mathbf{S}}^{(t)}$ and the cavity distribution stored in $\mathbf{P}^{(t)}$. The output of the GRU will be the input of the GNN module at the next EP iteration. The readout module $R$ in (30) computes the mean and variance estimation of each $x_{i}$. Two GNN layers perform (27)-(29) two times. More explanations are given in the following two subsections.
@@ VI-C Data-dependent Graph Convolution Coefficients
In equation (5), ChebNet uses conventional Chebyshev polynomial as graph filter, which uses fixed coefficients $w_{m}$. However, this architecture is less powerful and cannot perform well on testing data. Thus, in (28), we proposed to replace the fixed $w_{m}$ by data-dependent learnable attention scores $\text{ATT}(\mathbf{S}^{(0)})_{m}$. Training $\text{ATT}(\mathbf{S}^{(0)})$, which is a function of an instance $\{{\mathbf{y}},\mathbf{H},\sigma_{n}\}$, can improve the generalization ability of the graph convolution module. $\text{ATT}(\mathbf{S}^{(0)})$ first computes attention score $\alpha_{i}$ between the default coefficient vector ${\mathbf{1}}$ with size $M+1$ and each row of the key matrix $\mathbf{K}\in{\mathbb{R}}^{2N_{t}\times(M+1)}$,
$\alpha_{i}=\frac{\exp(\mathbf{K}_{i,:}{\mathbf{1}})}{\sum_{j}\exp(\mathbf{K}_{j,:}{\mathbf{1}})},\;\text{where }\mathbf{K}=\text{MLP}_{3}(\mathbf{S}^{(0)}).$ (31)
The $m$-th column of $\mathbf{K}$ is an embedding for the $m$-th coefficient. The output of $\text{ATT}(\mathbf{S}^{(0)})$ is a weighted average of each dimension of the coefficient embedding in rows of $\mathbf{K}$, where the attention score $\alpha_{i}$ serves as the normalized weight.
$\text{ATT}(\mathbf{S}^{(0)})=\left(\textstyle\sum_{i}\alpha_{i}\mathbf{K}_{i,:}\right)^{\mkern-1.5mu\mathsf{T}}.\textstyle$ (32)
@@ VII Computational Complexity Analysis
For GCEPNet, the $\text{MLP}_{1}$ in (27) has an input and an output layer of size $N_{u}$, and the $\text{MLP}_{3}$ in (31) has an input of size $2$ and an output of size $M+1$, then the total cost for the graph convolution in (21), (27), (28), (31) and (32) is $\mathcal{C}_{2}=N(N_{u}N_{h_{1}}+N_{h_{1}}N_{h_{2}}+N_{h_{2}}N_{u}+N_{h_{2}}M+M)+N^{2}N_{u}M$. When $M<433$, $\mathcal{C}_{2}$ has a smaller coefficient of $N^{2}$ than $\mathcal{C}_{1}$. In practice, we use $M=3$. Figure 2 demonstrates that, comparing with EP, GEPNet  [^2] Following the implementations in [16]: https://github.com/GNN-based-MIMO-Detection/GNN-based-MIMO-Detection does not scale with the problem size as a result of the inefficient GNN aggregation, while GCEPNet effectively resolves the bottleneck with the newly proposed graph convolution.
Figure 2: The total inference running time (in seconds) for $5\times 10^{4}$ samples with $N_{t}=N_{r}$ and the same number of EP iterations ($T=9$). Both GCEPNet and GEPNet use 2 GNN layers in each EP iteration. The implementation of GEPNet is from [16]. All methods are under CUDA acceleration.
@@ VIII Numerical Results
The training uses 850 epochs to optimize parameters in each network to minimize the average sample loss. There are 100 iterations per epoch and each iteration contains 100 training samples. Every training sample has a $\text{SNR}_{\text{dB}}$ uniformly sampled from $[25,50]$. In each epoch, the network is tested on a validation set containing 2000 samples for each $\text{SNR}_{\text{dB}}$ from $\{25,26,\cdots,50\}$. The training uses the Adam optimizer with a learning rate of $10^{-3}$, which is adjusted very 100 epochs according to the validation loss by the ReduceLROnPlateau scheduler in PyTorch. The network with the best mean SER on the validation set is selected for the testing stage. All networks are trained for each problem size and QAM configuration.
The testing uses $10^{5}$ samples for each $\text{SNR}_{\text{dB}}$. Figure 3 compares the average SER of GCEPNet, GEPNet [16] and EP [9] for $N_{t}=N_{r}=16$ and $N_{t}=N_{r}=32$ with 64-QAM. Both GCEPNet and GEPNet enhance the performance of EP by incorporating correlation information into the standard cavity estimation. While GCEPNet is significantly more efficient than GEPNet, it consistently outperforms GEPNet as well. Given that GEPNet achieves the best performance among existing sub-optimal detectors [17], GCEPNet is established as the new state-of-the-art (SOTA) method.